OpenStudy (anonymous):

Please help : The absoluyte value of y-2-4y=1

OpenStudy (liamschumm):

What is \[\sqrt{x^2}?\]

OpenStudy (anonymous):

isnt it 1 sorry im not so good at math

OpenStudy (tkhunny):

\(|y^{-2} - 4y|\)?? Who told you that? Answer the question, anyway.

OpenStudy (tkhunny):

You throw an x and there and somehow it turns into "1"? Why would that happen.? Try again.

OpenStudy (liamschumm):

It is actually |x|.

OpenStudy (liamschumm):

Which is the absolute value of x.

OpenStudy (anonymous):

ok got it

OpenStudy (anonymous):

so then how would i solve this?

OpenStudy (tkhunny):

We still don't know what to solve. Please write it more clearly. Use the vertical bar symbol for Absolute Value.

OpenStudy (anonymous):

Iy-2I-4y=1 (the abosolute value of y-2 )

OpenStudy (anonymous):



OpenStudy (tkhunny):

Ah! That makes so much more sense. You could have used parentheses in the first place to make it more clear on your first attempt. Iy-2I-4y=1 You MUST get this idea in your head: IF x > 0, the |x| = x. Absolute values don't do anything to positive numbers. |3| = 3 for example. Follow so far?

OpenStudy (tkhunny):

Follow or not, you need this, too... IF x < 0, the |x| = -x. Absolute values change the sign of negative numbers. |-2| = 2 for example. You have to get this stuck in your head.

OpenStudy (anonymous):

ok got it

OpenStudy (tkhunny):

To your problem... IF y - 2 > 0, we have an identical problem (y-2)-4y=1 Can you solve that? IF y - 2 < 0, we have an identical problem -(y-2)-4y=1 Can you solve that?